# Exploring the magnetic landscape of easily-exfoliable two-dimensional materials

### v1

@@ Abstract
Among metallic systems, we are particularly interested in half-metallic ferromagnetic systems that are interesting for spintronics devices because they can provide fully spin-polarized currents. 88 We thus search for systems where the Fermi energy crosses the energy bands for one spin channel while the other spin channel remains gapped. 89 Twelve novel half-metals are identified, whose electronic structure is reported in fig. 8. In several cases the bandwidth is sufficiently large to expect spin-polarized transport even in the presence of disorder. Remarkably, in systems like CuO2, Cr3O8, EuOI, EuOBr2, VO3, VS2O8, YbBr3, and YbCl3 a large gap is present above the metallic spin channel, which is potentially interesting in view of possible optoelectronic control of spin currents. In all these cases a correct description of the magnetic ground state is crucial to have half-metallic properties. For instance, in the case of FeAl2S4 it is important to have the two iron atoms with different oxidation states (+1 and +2) and thus different magnetic moments of 2.2 and 3.2 . We also stress that, in general, the half-metallic state can be fragile, especially if it is protected by symmetry-induced degeneracies. 90 Further tests to include the effect of spin-orbit coupling or structural distortions would thus be needed to fully validate predictions, although the procedure outlined above is already sufficient to discard speculative half-metals such as iron dihalides FeX2 (with X = Cl, Br, I). These systems have been predicted to be half-metals, 91; 92 although experimental evidence 93; 94; 95; 96; 97suggests that they have a finite gap. We find them to be indeed insulating, although the mechanism to break the symmetry-induced degeneracy arises from an incorrect AFM configuration, as we don’t include relativistic effects or distortions.
I Conclusions
@@
In conclusion, we investigate the magnetic ground states of easily exfoliable monolayers from the MC2D database. Starting from 3077 easily and potentially exfoliable monolayers identified in Refs. 14; 13, an initial screening is carried out on 877 easily-exfoliable systems containing up to 12 atoms per cell using high-throughput density functional theory with the conventional PBE exchange-correlation functional. Testing various magnetic configurations by changing the initial magnetization, 228 magnetic monolayers are identified as potentially magnetic systems. Subsequently, Hubbard corrections are introduced for these systems, with the Hubbard parameter determined self-consistently. The magnetic energy landscape is then thoroughly explored by applying constraints to the occupation matrices of the or orbitals, enabling the identification of the global energy minimum. Through this process, we identify 194 magnetic monolayers, comprising 109 ferromagnetic, 83 antiferromagnetic, and 2 altermagnetic systems. To estimate critical temperatures and highlight promising candidates, we employ a proxy based on the energy difference between ferromagnetic and antiferromagnetic states. This method is applied to the 137 monolayers that exhibit an electronic band gap. Among the remaining 57 metallic systems, we identify 12 novel half-metals with strong potential for spintronics applications.
@@ Methods
In this work, all calculations are collinear and performed using the Quantum ESPRESSO (QE) distribution. 98; 99; 100 Calculations are performed within the
@@
spin-polarized generalized-gradient approximation (GGA) for exchange-correlation functional using the Perdew-Burke-Ernzerhof (PBE) prescription. The pseudopotentials and the wavefunction kinetic-energy cutoffs are taken from a standard solid-state pseudopotentials (SSSP Efficiency) library. 101; 102; 103; 104; 105 We use the most recent versions of the SSSP library available at the time for our calculations. Specifically, version 1.1.2 is used for the PBE calculations, and version 1.3.0 is utilized for the PBE+ calculations. The -points density of and -points mesh density of have been set for single self-consistent (scf) and density functional perturbation theory (DFPT) calculations, respectively. In this work, a system (an atom) is magnetic if the absolute total magnetization per cell (of the atom) is larger than 0.1. The data used to produce some of the results of this paper are given in the materials cloud archive. 106
@@
In this work we consider systems from MC2D, 58; 59; 14 which is created from a collection of 3077 exfoliable monolayer that were identified from prototypes of the Inorganic Crystal Structure Database (ICSD), 107; 108 the Crystallography Open Database (COD), 109 and Materials Platform for Data Science (MPDS) 110 databases. We identify and characterize magnetic monolayers in MC2D in two main steps, as shown in fig. 1: In the first step (Chronos step), using PBE calculations, potentially magnetic monolayers are identified, and in the second step (RomeoDFT step), after including Hubbard corrections and exploring the energy landscape, the magnetic ground state of these materials is found. An initial screening was performed using an automated AiiDA 60; 61 workflow (Chronos) on 877 easily exfoliable materials (i.e. materials with a binding energy per unit of area lower or equal to 30 meV/Å2 when computed with the vdW-DF2-c09 functional or lower or equal to 35 meV/Å2 when computed with the rvv10) with up to 12 atoms per unit cell, using PBE calculations, and 228 magnetic monolayers are identified by following procedure: the geometry of the system is optimized in ferromagnetic and non-magnetic orderings. If the ferromagnetic ordering shows lower energy than the non-magnetic case, the system is considered to be potentially magnetic, and other antiferromagnetic orderings (making sure the supercell size is now to accommodate two magnetic atoms) and five random configurations are created. The system is optimized starting from different starting magnetization defined by spin density ( where describes the density and I is the atom index). As stated before this simplistic approach is effective in capturing the correct magnetic ground state in the absence of Hubbard correction (see Supporting Information). However, it is important to note that there are still local minima within the energy landscape that are accessible through the manipulations of the orbital occupations.
@@
where EU is the Hubbard corrected energy, is the Lagrange multiplier, representing the strength of the energy penalty associated with a deviation from the target occupation matrices, is the atomic site index, and are the magnetic quantum numbers associated with a specific angular momentum and is the on-site Hubbard parameters. are the generalized atomic occupation matrices which are computed by projecting the Kohn-Sham (KS) wavefunctions on atomic orbitals as: , where are the occupations of KS states. The general formulation of DFT+ is discussed in Refs. 30; 37; 35; 33. We note that in this case the magnetization is defined from the occupation matrix of spin up and down: . After imposing the constraint for a fixed number of self-consistent cycles, the constraint is released and the system is allowed to evolve freely to the closest self-consistent energy minimum. More information about RomeoDFT is provided in Ref. 48. This implementation is made in Quantum ESPRESSO and the process is automatized in Julia and can be found on GitHub. We use RomeoDFT workflow to determine the ground state at the PBE+ level for 194 magnetic monolayers, identified by Chronos with a filter of up to 12 atoms per primitive cell.
In order to calculate the parameter, we use DFPT 35; 33 as implemented in the HP code, 34; 31; 37 which is part of QE. The Hubbard parameter is computed in three iterative steps. Within each iteration the value is updated as the average values of the initial and calculated values in the previous step. 111Only for special case of CoO2, because the ground state is highly sensitive to the choice of , in order to avoid falling into the wrong states we do not update the value as the average of the previous step, instead, in each step we use the same calculated value of the from the previous step. Our results show that with this approach, the convergence threshold of for 87 of the systems is less than 0.5 eV. If the convergence of 0.5 eV has not been reached in three steps, we do another final step. After the calculation of the , the geometry and the Hubbard are fixed, and we explore the energy landscape to find the global minimum. For the systems with fully occupied orbitals we use a zero value for the because the effects of the Hubbard correction at integer occupations are negligible. 31
@@
where is the Boltzmann constant and . We calculated the exchange parameters using two approaches. First, we employed the method, which maps this total energy difference onto a nearest-neighbor-only model Hamiltonian and the results are shown in fig. 6. Here, the exchange coupling is assumed to be present between equivalent nearest neighbors () and given by , where is the energy difference between self-consistent ferromagnetic and antiferromagnetic spin configurations and the number of magnetic atoms in the unit cell. To account for interactions beyond the first nearest neighbors, we employ the four-state method. In this approach, a large supercell is considered (), and each exchange interaction is calculated by computing the total energy when the pair of spins in sites and are in four distinct collinear spin configurations while keeping all other spins ferromagnetically ordered. The four configurations are up-up, down-up, up-down, and down-down concerning the quantization axis. The exchange parameter is then calculated by: 84
@@
Finally, to go beyond the mean-field theory and obtain a more precise estimate of the paramagnetic critical temperatures, we use the Monte Carlo Metropolis algorithm as implemented in the Vampire package 112; 113 for a few selected materials. Here, we use periodic boundaries in 2D, an equilibration time of 5.000 Monte Carlo steps, and statistical averaging also over 5.000 steps. The critical temperatures are then obtained by fitting the Monte Carlo magnetization curves to the mean-field formula , where and the critical exponent are also fitting parameters
@@
It is important to note the approximations involved in our results. Besides assuming a magnetism driven by localized magnetic moments, the mapping of the onto the Heisenberg model considers only nearest neighbors coupling. This can be regarded as a mean-field approximation. Take the case of MoCl3, where the Mo atoms go from a honeycomb lattice arrangement, with three equidistant nearest-neighbor sites, to a triangular lattice of dimers where the distance between Mo sites in the dimer is 1.06 Å smaller than between dimers (see Supporting Information).
@@
Anisotropy is also neglected in our calculations, and all calculations are done within the non-relativistic limit. While anisotropy could play a significant role in 2D materials, our MC simulations on a model honeycomb lattice considering only the nearest neighbors indicate that is not affected significantly by a typical magnetocrystalline anisotropy. This aligns with recent research, that shows that even for isotropic 2D systems in lab setups short-range interactions can be large enough to stabilize the magnetic order at a finite temperature due to finite effects. 85
Acknowledgements.
II acknowledgments
@@ Acknowledgements.
We thank Guoyuan Liu for fruitful discussions. We acknowledge support from the Swiss National Science Foundation (SNSF), through grant 200021-179138, and its National Centre of Competence in Research (NCCR) MARVEL (grant number 205602). MG acknowledges financial support from Ministero Italiano dell’Università e della Ricerca through the PNRR project ”Ecosystem for Sustainable Transition in Emilia-Romagna” (ECS_00000033-ECOSISTER) and through the PRIN2022 project SECSY (CUP E53D23001700006), both funded by the European Union – NextGenerationEU. Computer time was provided by the Swiss National Supercomputing Centre (CSCS) under project No. s1073. The authors declare no competing financial interests.

### v2

@@ Abstract
Among metallic systems, we are particularly interested in half-metallic ferromagnetic systems that are interesting for spintronicss devices because they can provide fully spin-polarized currents. 89 We thus search for systems where the Fermi energy crosses the energy bands for one spin channel while the other spin channel remains gapped. 90 Twelve novel half-metals are identified, whose electronic structure is reported in Figure 8. In several cases the bandwidth is sufficiently large to expect spin-polarized transport even in the presence of disorder. Remarkably, in systems like CuO2, Cr3O8, EuOI, EuOBr2, VO3, VS2O8, YbBr3, and YbCl3 a large gap is present above the metallic spin channel, which is potentially interesting in view of possible optoelectronic control of spin currents. In all these cases a correct description of the magnetic ground state is crucial to have half-metallic properties. For instance, in the case of FeAl2S4 it is important to have the two iron atoms with different oxidation states (+1 and +2) and thus different magnetic moments of 2.2 and 3.2 . We also stress that, in general, the half-metallic state can be fragile, especially if it is protected by symmetry-induced degeneracies. 91 Further tests to include the effect of spin-orbit coupling or structural distortions would thus be needed to fully validate predictions, although the procedure outlined above is already sufficient to discard speculative half-metals such as iron dihalides FeX2 (with X = Cl, Br, I). These systems have been predicted to be half-metals, 92, 93 although experimental evidence 94, 95, 96, 97, 98suggests that they have a finite gap. We find them to be indeed insulating, although the mechanism to break the symmetry-induced degeneracy arises from an incorrect AFM configuration, as we do not include relativistic effects or distortions.
1 Conclusions
@@
In conclusion, we investigate the magnetic ground states of easily exfoliable monolayers from the MC2D database. Starting from 3077 easily and potentially exfoliable monolayers identified in Refs. 14, 13, an initial screening is carried out on 877 easily-exfoliable systems containing up to 12 atoms per cell using high-throughput density functional theory with the conventional PBE exchange-correlation functional. Testing various magnetic configurations by changing the initial magnetization, 228 magnetic monolayers are identified as potentially magnetic systems. Subsequently, Hubbard corrections are introduced for these systems, with the Hubbard parameter determined self-consistently. The magnetic energy landscape is then thoroughly explored by applying constraints to the occupation matrices of the or orbitals, enabling the identification of the global energy minimum. Through this process, we identify 194 magnetic monolayers, comprising 109 ferromagnetic, 83 antiferromagnetic, and 2 altermagnetic systems. To estimate critical temperatures and highlight promising candidates, we employ a proxy based on the energy difference between ferromagnetic and antiferromagnetic states. This method is applied to the 137 monolayers that exhibit an electronic band gap. Among the remaining 57 metallic systems, we identify 12 novel half-metals with strong potential for spintronics applications.
@@ Methods
In this work, all calculations are collinear and performed using the Quantum ESPRESSO (QE) distribution. 99, 100, 101 Calculations are performed within the
@@
spin-polarized generalized-gradient approximation (GGA) for exchange-correlation functional using the Perdew-Burke-Ernzerhof (PBE) prescription. The pseudopotentials and the wavefunction kinetic-energy cutoffs are taken from a standard solid-state pseudopotentials (SSSP Efficiency) library. 102, 103, 104, 105, 106 We use the most recent versions of the SSSP library available at the time for our calculations. Specifically, version 1.1.2 is used for the PBE calculations, and version 1.3.0 is utilized for the PBE+ calculations. The -points density of and -points mesh density of have been set for single self-consistent (scf) and density functional perturbation theory (DFPT) calculations, respectively. In this work, a system (an atom) is magnetic if the absolute total magnetization per cell (of the atom) is larger than 0.1. The data used to produce some of the results of this paper are given in the materials cloud archive. 107
@@
In this work we consider systems from MC2D, 59, 60, 14 which is created from a collection of 3077 exfoliable monolayer that were identified from prototypes of the Inorganic Crystal Structure Database (ICSD), 108, 109 the Crystallography Open Database (COD), 110 and Materials Platform for Data Science (MPDS) 111 databases. We identify and characterize magnetic monolayers in MC2D in two main steps, as shown in Figure 1: in the first step (Chronos step), using PBE calculations, potentially magnetic monolayers are identified, and in the second step (RomeoDFT step), after including Hubbard corrections and exploring the energy landscape, the magnetic ground state of these materials is found. An initial screening was performed using an automated AiiDA 61, 62 workflow (Chronos) on 877 easily exfoliable materials (i.e. materials with a binding energy per unit of area lower or equal to 30 meV/Å2 when computed with the vdW-DF2-c09 functional or lower or equal to 35 meV/Å2 when computed with the rvv10) with up to 12 atoms per unit cell, using PBE calculations, and 228 magnetic monolayers are identified by the following procedure: the geometry of the system is optimized in ferromagnetic and non-magnetic orderings. If the ferromagnetic ordering shows lower energy than the non-magnetic case, the system is considered to be potentially magnetic, and other antiferromagnetic orderings (making sure the supercell size is now to accommodate two magnetic atoms) and five random configurations are created. The system is optimized starting from different starting magnetization defined by spin density ( where describes the density and I is the atom index). As stated before this simplistic approach is effective in capturing the correct magnetic ground state in the absence of Hubbard correction (see Supporting Information). However, it is important to note that there are still local minima within the energy landscape that are accessible through the manipulations of the orbital occupations.
@@
where EU is the Hubbard corrected energy, is the Lagrange multiplier, representing the strength of the energy penalty associated with a deviation from the target occupation matrices, is the atomic site index, and are the magnetic quantum numbers associated with a specific angular momentum and is the on-site Hubbard parameters. are the generalized atomic occupation matrices which are computed by projecting the Kohn-Sham (KS) wavefunctions on atomic orbitals as: , where are the occupations of KS states. The general formulation of DFT+ is discussed in Refs. 30, 37, 35, 33. We note that in this case the magnetization is defined from the occupation matrix of spin up and down: . After imposing the constraint for a fixed number of self-consistent cycles, the constraint is released and the system is allowed to evolve freely to the closest self-consistent energy minimum. More information about RomeoDFT is provided in Ref. 48. This implementation is made in Quantum ESPRESSO and the process is automatized in Julia and can be found on GitHub. We use RomeoDFT workflow to determine the ground state at the PBE+ level for 194 magnetic monolayers, identified by Chronos with a filter of up to 12 atoms per primitive cell.
In order to calculate the parameter, we use DFPT 35, 33 as implemented in the HP code, 34, 31, 37 which is part of QE. The Hubbard parameter is computed in three iterative steps. Within each iteration the value is updated as the average values of the initial and calculated values in the previous step. 111Only for special case of CoO2, because the ground state is highly sensitive to the choice of , in order to avoid falling into the wrong states we do not update the value as the average of the previous step, instead, in each step we use the same calculated value of the from the previous step. Our results show that with this approach, the convergence threshold of for 87 of the systems is less than 0.5 eV. If the convergence of 0.5 eV has not been reached in three steps, we do another final step. After the calculation of the , the geometry and the Hubbard are fixed, and we explore the energy landscape to find the global minimum. For the systems with fully occupied orbitals we use a zero value for the because the effects of the Hubbard correction at integer occupations are negligible. 31
@@
where is the Boltzmann constant and . We calculated the exchange parameters using two approaches. First, we employed the method, which maps this total energy difference onto a nearest-neighbor-only model Hamiltonian and the results are shown in Figure 6. Here, the exchange coupling is assumed to be present between equivalent nearest neighbors () and given by , where is the energy difference between self-consistent ferromagnetic and antiferromagnetic spin configurations and the number of magnetic atoms in the unit cell. To account for interactions beyond the first nearest neighbors, we employ the four-state method. In this approach, a large supercell is considered (), and each exchange interaction is calculated by computing the total energy when the pair of spins in sites and are in four distinct collinear spin configurations while keeping all other spins ferromagnetically ordered. The four configurations are up-up, down-up, up-down, and down-down concerning the quantization axis. The exchange parameter is then calculated by: 85
@@
Finally, to go beyond the mean-field theory and obtain a more precise estimate of the paramagnetic critical temperatures, we use the Monte Carlo Metropolis algorithm as implemented in the Vampire package 112, 113 for a few selected materials. Here, we use periodic boundaries in 2D, an equilibration time of 5.000 Monte Carlo steps, and statistical averaging also over 5.000 steps. The critical temperatures are then obtained by fitting the Monte Carlo magnetization curves to the mean-field formula , where and the critical exponent are also fitting parameters
@@
It is important to note the approximations involved in our results. Besides assuming a magnetism driven by localized magnetic moments, the mapping of the onto the Heisenberg model considers only nearest neighbors coupling. This can be regarded as a mean-field approximation. For instance, in MoCl3, the Mo atoms transition from a honeycomb lattice, where each site has three equidistant nearest neighbors, to a triangular lattice of dimers. In this configuration, the Mo–Mo distance within a dimer is 1.06 Å shorter than the distance between dimers (see Supporting Information).
@@
Anisotropy is also neglected in our calculations, and all calculations are done within the non-relativistic limit. While anisotropy could play a significant role in 2D materials, our MC simulations on a model honeycomb lattice considering only the nearest neighbors indicate that is not affected significantly by a typical magnetocrystalline anisotropy. This aligns with recent research, that shows that even for isotropic 2D systems in lab setups short-range interactions can be large enough to stabilize the magnetic order at a finite temperature due to finite effects. 86
@@
We thank Giovanni Pizzi and Andrea Cepelotti for their crucial roles in helping with scientific discussions and workflow development. We also Guoyuan Liu for fruitful discussions. We acknowledge support from the Swiss National Science Foundation (SNSF), through grant 200021-179138, and its National Centre of Competence in Research (NCCR) MARVEL (grant number 205602). MG acknowledges financial support from Ministero Italiano dell’Università e della Ricerca through the PNRR project ”Ecosystem for Sustainable Transition in Emilia-Romagna” (ECS_00000033-ECOSISTER) and through the PRIN2022 project SECSY (CUP E53D23001700006), both funded by the European Union – NextGenerationEU. Computer time was provided by the Swiss National Supercomputing Centre (CSCS) under project No. s1073. The authors declare no competing financial interests.